please help me to solve:
Find a ratio for each of the following, and write it as a ratio of intergers in lowest terms.
24) Coffee Drinkers: Among 100 coffee drinkers, 36 said that they preferred their coffee black and the rest did not prefer their coffee black. Find the ratio of those who prefer black to those who prefer non black coffee.
Thank you.

Answer by malakumar_kos@yahoo.com(315)

Find a ratio for each of the following, and write it as a ratio of intergers in lowest terms.
24) Coffee Drinkers: Among 100 coffee drinkers, 36 said that they preferred their coffee black and the rest did not prefer their coffee black. Find the ratio of those who prefer black to those who prefer non black coffee.

Total number of coffee drinkers = 100
number preferring black coffee = 36
number not preferring black coffee = 64 (100-36=64)
Ratio of integers in lowest terms = 36/64
= 9/16